# ScalingNoise: Scaling Inference-Time Search for Generating Infinite Videos

### v2

@@ Abstract
Video diffusion models (VDMs) facilitate the generation of high-quality videos, with current research predominantly concentrated on scaling efforts during training through improvements in data quality, computational resources, and model complexity. However, inference-time scaling has received less attention, with most approaches restricting models to a single generation attempt. Recent studies have uncovered the existence of “golden noises” that can enhance video quality during generation. Building on this, we find that guiding the scaling inference-time search of VDMs to identify better noise candidates not only evaluates the quality of the frames generated in the current step but also preserves the high-level object features by referencing the anchor frame from previous multi-chunks, thereby delivering long-term value. Our analysis reveals that diffusion models inherently possess flexible adjustments of computation by varying denoising steps, and even a one-step denoising approach, when guided by a reward signal, yields significant long-term benefits. Based on the observation, we propose ScalingNoise, a plug-and-play inference-time search strategy that identifies golden initial noises for the diffusion sampling process to improve global content consistency and visual diversity. Specifically, we perform one-step denoising to convert initial noises into a clip and subsequently evaluate its long-term value, leveraging a reward model anchored by previously generated content. Moreover, to preserve diversity, we sample candidates from a tilted noise distribution that up-weights promising noises. In this way, ScalingNoise significantly reduces noise-induced errors, ensuring more coherent and spatiotemporally consistent video generation. Extensive experiments on benchmark datasets demonstrate that the proposed ScalingNoise effectively improves both content fidelity and subject consistency for resource-constrained long video generation.
@@
11footnotetext: Equal Contribution. Corresponding Authors.
1 Introduction
Long video generation has a significant impact on various applications, including film production, game development, and artistic creation [39, 87, 94]. Compared to image generation [80, 41, 16], video generation demands significantly greater data scale and computational resources due to the high-dimensional nature of video data. This necessitates a trade-off between limited resources and model performance for Video Diffusion Models (VDMs) [35, 48, 82, 65].
Recent VDMs typically adopt two main methods: one is the chunked autoregressive strategy [26, 23, 72, 43, 4], which predicts several frames in parallel conditioned on a few preceding ones, consequently reducing the computational burden, and “diagonal denoising” from FIFO-diffusion [31, 11], which re-plans the time schedule and maintains a queue with progressively increasing noise levels for denoising. However, video generation is induced by both the diffusion strategies and the noise. Variations in the noises can lead to substantial changes in the synthesized video, as even minor alterations in the noise input can dramatically influence the output [89, 51, 115]. This sensitivity underscores that noises affect both the overall content and the subject consistency of long video generation.
@@ 1 Introduction
Given limited computational resources, our strategy selects from a finite pool of candidate noises. Moreover, the quality of candidate noises constitutes a critical factor, complementing the robust reward model that provides long-term supervisory signals. To ensure the candidate pool comprises noise that enhances video consistency, we construct it by sampling from a skewed distribution. Specifically, the weight of high-reward samples is increased, while samples are still drawn from the standard normal distribution to preserve diversity. Through this iterative search process, we significantly reduce accumulated errors and avoid inconsistencies arising from the randomness of initial noises.
On multiple benchmarks, we verify the effectiveness of our method. Our main contributions are as follows:
  • We propose a plug-and-play inference-time scaling strategy for long video generation by incorporating long-term supervision signals into the reward process.
  • We introduce a one-step denoising approach that transforms the evaluation of initial noises into the evaluation of a clearer image without extra computational overhead.
  • Extensive experiments demonstrate that our proposed ScalingNoise can be effectively applied to various video diffusion models. It consistently improves the inference quality of generated videos.
2 Related Work
Long viedo generation. Video generation has advanced significantly [95, 59, 3, 113, 69, 112], yet producing high-quality long videos remains challenging due to the scarcity of such data and the high computational resources required [10, 3, 77]. This limits training models for direct long video generation, leading to the widespread use of autoregressive approaches built on pre-trained models [15, 91]. Current solutions fall into two categories: training-based [76, 38, 109, 58] and training-free methods [110, 5, 107, 102]. Training-based methods [84, 19] like NUWA-XL [96] use a divide-and-conquer strategy to generate long videos layer by layer, while FDM [22] and SEINE [12] combine frame interpolation and prediction. Other approaches, such as LDM [4], MCVD [72], Latent-Shift [2], and StreamingT2V [24], incorporate short-term and long-term information as additional inputs. Despite their success, these methods demand high-quality long video data and substantial computation. Training-free methods address these challenges. Gen-L-Video [74] and Freenoise [53] use a chunk-by-chunk approach, linking segments with final frames, but this risks degradation and inconsistency. Freelong [42] blends global and local data via high-low frequency decoupling, while FreeInit [82] refines initial noises for better consistency. FIFO-Diffusion [31] introduces a novel paradigm, reorganizing denoising with a noise-level queue, dequeuing clear frames, enqueuing noise, for efficient, flexible long video generation without training. In this work, we propose a plug-and-play inference-time strategy that can improve the consistency of videos based on these long video generation methods.
Inference-time Search. A variety of inference-time search strategies have been proven crucial in long context generation within the field of LLMs [99, 29, 34, 66, 92, 40]. The advent of DeepSeek-R1 [14] has further advanced inference-time search. By applying various search techniques in the language space, such as controlled decoding [7, 90, 18, 100], best of N [36, 33], and Monte Carlo tree search [105, 67, 73, 79], LLMs achieve better step-level responses, thus enhancing performance. During inference-time search, leveraging a good process reward model (PRM) [70, 13, 27, 78] is essential to determine the quality of the responses. Ma et al. [45] proposed using supervised verifiers as a signal to guide generating trajectory within Diffusion Models (DMs), but did not investigate its impact on video generation inference-time search. Furthermore, some work has preliminarily explored inference-time search in VDMs [69, 85, 104, 68], however, there is still a lack of investigation into the inference-time scaling law and long-term signals in the process of long video generation. In this work, we explore the effectiveness of scaling inference-time budget utilizing beam search to enhance the consistency of generated long videos.
@@ 3.1 Preliminaries: Video Denoising Approach
Long Video Generation. We introduce the formulation of VDM for generating long videos. There are two approaches to create a long video: the chunk-by-chunk method and the FIFO diffusion method. Both paradigms can be seamlessly integrated into the subsequent framework. Consider extending a pre-trained model that generates fixed-length videos into a long video generation model with a distribution of $p_{\theta}$, based on the two approaches mentioned above. This model processes an input consisting of both prompt and image to generate a video $\mathbf{v}=[v_{1},v_{2},\dots,v_{N}]$, where $\mathbf{v}$ consists of N step-level responses. Each step-level response $v_{i}$ is a frame or video clip of the long video, treated as a sample drawn from a conditional probability distribution:
$v_{i}=p_{\theta}(v_{i}|\mathbf{z}_{<i},c),\quad i=1,2,\dots,T,$ (1)
where $c$ can be denoted as a single prompt or a prompt, image pair and $\mathbf{z}_{<i}=[v_{1},v_{2},\dots,v_{i}]$ is the concatenated video. In the following, we provide the specific step-level formulations for these two paradigms, respectively:
@@
    where $\Psi$ and $\bm{\epsilon}_{\theta}$ denote a sampler such as DDIM and a noise predict network, respectively.
  • FIFO-Diffusion: Different from the aforementioned process, FIFO-Diffusion [31] introduces a diagonal denoising paradigm [56, 9] by rescheduling noise steps during the inference process. FIFO-Diffusion achieves autoregressive generation by maintaining a queue where the noise level increases step by step. We define the queue as $Q=\left\{v_{i,t}\right\}_{t=1}^{T}$, where $t$ denotes the noise level, and $i$ indicates the $i$-th frame in the queue. In this paradigm, $i$ is equal to $t$, and $v_{i}$ denotes a frame. The length of $T$ is $M\times P$ where $P$ denotes the partition of the queue. The procedure of FIFO step can be described as follows:
@@
$\begin{split}\small v_{\tau_{t-1}}&=\text{DDIM}(v_{\tau_{t}})\\ &=\alpha_{\tau_{t-1}}\left(\frac{v_{\tau_{t}}-\sigma_{\tau_{t}}\bm{\epsilon}_{\theta}(v_{\tau_{t}},\tau_{t})}{\alpha_{\tau_{t}}}\right)+\sigma_{\tau_{t-1}}\bm{\epsilon}_{\theta}(v_{\tau_{t}},\tau_{t}),\end{split}$ (4)
@@ 3.2 Formulation of Video Generation Inference
The long video generation framework can be formulated as a Markov Decision Process (MDP) problem defined by the tuple $(\mathcal{S},\mathcal{A},\mathcal{R})$. $\mathcal{S}$ is the state space. Each state is defined as a combination of the generated video and the condition. The initial state $\bm{s}_{0}$ only corresponds to the input prompt $x$ and guided image $I$. $\bm{S}_{i}$ is the combination of the currently generated videos. $\mathcal{A}$ denotes the action space, where each action encompasses a two-part process: sampling an initial noise from a tilted distribution, followed by denoising the current video clip based on the noise. We also have the reward function $\mathcal{R}$ to evaluate the reward of each action, which is also known as the process reward model (PRM) in LLMs. With this MDP modeling, we can search for additional states by increasing the inference-time compute, thereby obtaining a better VDM response $\mathbf{v}$. Specifically, we can take different actions in each state, continuously explore, and then make choices based on the reward model to achieve a better state. The general formulation of the selection process is:

### v3

@@ Abstract
Video diffusion models (VDMs) facilitate the generation of high-quality videos, with current research predominantly concentrated on scaling efforts during training through improvements in data quality, computational resources, and model complexity. However, inference-time scaling has received less attention, with most approaches restricting models to a single generation attempt. Recent studies have uncovered the existence of “golden noises" that can enhance video quality during generation. Building on this, we find that guiding the scaling inference-time search of VDMs to identify better noise candidates not only evaluates the quality of the frames generated in the current step but also preserves the high-level object features by referencing the anchor frame from previous multi-chunks, thereby delivering long-term value. Our analysis reveals that diffusion models inherently possess flexible adjustments of computation by varying denoising steps, and even a one-step denoising approach, when guided by a reward signal, yields significant long-term benefits. Based on the observation, we propose ScalingNoise, a plug-and-play inference-time search strategy that identifies golden initial noises for the diffusion sampling process to improve global content consistency and visual diversity. Specifically, we perform one-step denoising to convert initial noises into a clip and subsequently evaluate its long-term value, leveraging a reward model anchored by previously generated content. Moreover, to preserve diversity, we sample candidates from a tilted noise distribution that up-weights promising noises. In this way, ScalingNoise significantly reduces noise-induced errors, ensuring spatiotemporal coherence in video generation. Extensive experiments on benchmark datasets demonstrate that ScalingNoise effectively improves both content fidelity and subject consistency for resource-constrained long video generation.
1 Introduction
Long video generation has a significant impact on various applications, including film production, game development, and artistic creation [43, 93, 100]. Compared to image generation [86, 44, 16], video generation demands significantly greater data scale and computational resources due to the high-dimensional nature of video. This necessitates a trade-off between limited resources and model performance for Video Diffusion Models (VDMs) [37, 51, 88, 70].
11footnotetext: Equal Contribution. Corresponding Authors.
Recent VDMs typically adopt two main methods: one is the chunked autoregressive strategy [26, 23, 78, 46, 4], which predicts several frames in parallel conditioned on a few preceding ones, consequently reducing the computational burden, and “diagonal denoising” from FIFO-diffusion [33, 11], which re-plans the time schedule and maintains a queue with progressively increasing noise levels for denoising. However, video generation is induced by both the diffusion strategies and the noise. Variations in the noises can lead to substantial changes in the synthesized video, as even minor alterations in the noise input can dramatically influence the output [95, 56, 121]. This sensitivity underscores that noises affect both the overall content and the subject consistency of video generation.
@@ 1 Introduction
Given limited computational resources, our strategy selects from a finite pool of candidate noises. Moreover, the quality of candidate noises constitutes a critical factor, complementing the robust reward model that provides long-term supervisory signals. To ensure the candidate pool comprises noise that enhances video consistency, we construct it by sampling from a tilted distribution. Specifically, the weight of high-reward samples is increased, while samples are still drawn from the standard normal distribution to preserve diversity. Through this iterative search process, we significantly reduce accumulated errors and avoid inconsistencies arising from the randomness of initial noises.
On multiple benchmarks, we verify the effectiveness of our method. Our main contributions are: (i) We propose a plug-and-play inference-time scaling strategy for long video generation by incorporating long-term supervision signals into the reward process. (ii) We introduce a one-step denoising approach that transforms the evaluation of initial noises into the evaluation of a clearer image without extra computational overhead. (iii) Extensive experiments demonstrate that our proposed ScalingNoise can be effectively applied to various video diffusion models, improving the quality of generated videos.
@@ 2.1 Preliminaries: Video Denoising Approach
Long Video Generation. We introduce the formulation of VDM for generating long videos. There are two approaches to create long videos: the Chunk-by-Chunk method and the FIFO diffusion method. In the following, we provide the specific formulations for these two paradigms, respectively:
@@
    where $\Psi$ and $\bm{\epsilon}_{\theta}$ denote a sampler such as DDIM and a noise predict network, respectively, and $c$ can be denoted as a single prompt or a prompt, image pair.
  • FIFO-Diffusion: Different from the aforementioned process, FIFO-Diffusion [33] introduces a diagonal denoising paradigm [61, 9] by rescheduling noise steps. It achieves autoregressive generation by maintaining a queue where the noise level increases step by step. We define the queue as $Q=\left\{v_{i,t}\right\}_{t=1}^{T}$, where $t$ denotes the noise level, and $i$ indicates the $i$-th frame in the queue. In this paradigm, $i$ is equal to $t$, and $v_{i}$ denotes a frame. The length of $T$ is $M\times P$ where $P$ denotes the partition of the queue. The procedure of FIFO step can be described as follows:
@@
$v_{\tau_{t-1}}=\text{DDIM}(v_{\tau_{t}})=\alpha_{\tau_{t-1}}\left(\frac{v_{\tau_{t}}-\sigma_{\tau_{t}}\bm{\epsilon}_{\theta}(v_{\tau_{t}},\tau_{t})}{\alpha_{\tau_{t}}}\right)+\sigma_{\tau_{t-1}}\bm{\epsilon}_{\theta}(v_{\tau_{t}},\tau_{t}),$ (3)
@@ 2.2 Formulation of Video Generation Inference
Long video generation paradigms can be seamlessly integrated into the subsequent framework. Consider extending a pre-trained model that generates fixed-length videos into a long video generation model with a distribution of $p_{\theta}$. This model processes an input to generate a video $\mathbf{v}=[v_{1},v_{2},\dots,v_{N}]$, where $\mathbf{v}$ consists of N step-level responses. Each step-level response $v_{i}$ is a video clip of the long video, treated as a sample drawn from a conditional probability distribution:
$v_{i}=p_{\theta}(v_{i}|\mathbf{v}_{<i},c),\quad i=1,2,\dots,T,$ (4)
where $\mathbf{v}_{<i}=[v_{1},v_{2},\dots,v_{i}]$ denotes the concatenated video. Moreover, this framework can be formulated as a Markov Decision Process (MDP) problem defined by the tuple $(\mathcal{S},\mathcal{A},\mathcal{R})$. $\mathcal{S}$ is the state space. Each state is defined as a combination of the generated video and the condition. The initial state $\bm{s}_{0}$ only corresponds to the input. $\bm{s}_{i}$ is the combination of the currently generated videos. $\mathcal{A}$ denotes the action space, where each action encompasses a two-part process: sampling an initial noise from a tilted distribution, followed by denoising the current video clip based on the noise. We also have the reward function $\mathcal{R}$ to evaluate the reward of each action, which is also known as the process reward model (PRM) in LLMs. With this MDP modeling, we can search for additional states by increasing the inference-time compute, thereby obtaining a better VDM response $\mathbf{v}$. Specifically, we can take different actions in each state, continuously explore, and then make choices based on the reward model to achieve a better state. The general formulation of the selection process is: